The probability that a university graduate will be offered no jobs within the a month of graduation is estimated to be 10%. The probability of receiving one, two and three job offers has similarly been to be 38%,28% and 24%, and respectively. Determine the probability :
A. P(A graduate is offered fewer than two jobs)
B. P(A graduate is offered more than one job)

Answer by ikleyn(52777)

.
The probability that a university graduate will be offered no jobs within the a month of graduation
is estimated to be 10%. The probability of receiving one, two and three job offers has similarly been
to be 38%,28% and 24%, and respectively. Determine the probability :
A. P(A graduate is offered fewer than two jobs)
B. P(A graduate is offered more than one job)
~~~~~~~~~~~~~~

(A)  P(a graduate is offered fewer than two jobs)) = P(0) + P(1) = 0.1 + 0.38 = 0.48.    ANSWER


(B)  P(more than one job) = COMPLEMENTARY probability to not gettting any offer = 1 - 0.1 = 0.9.    ANSWER

Solved.